# The Curious Case of In-Training Compression of State Space Models

### v2

@@ 1 Introduction
State Space Models (SSMs) Gu et al. (2021); Hasani et al. (2022); Smith et al. (2022); Orvieto et al. (2023); Rusch & Rus (2025) have recently emerged as a powerful alternative to established sequence models such as Recurrent Neural Networks (RNNs) and Transformers. They combine the parallelizable training efficiency of scaled dot-product attention Vaswani (2017) with the computational and memory advantages of RNNs, enabling strong performance across large-scale language, vision, and audio modeling tasks Gu & Dao (2024); Goel et al. (2022); Nguyen et al. (2022).
Despite their efficient structure and recent progress in hardware-aware implementations, current SSMs remain computationally-intensive. While both memory and runtime scale with sequence length, the size of the SSM state further amplifies these costs. Reducing the state dimension therefore provides an effective strategy to simultaneously reduce memory usage and runtime. This can be achieved by leveraging techniques from structured compression. However, most existing approaches are commonly applied post-training: a large model is trained to completion and only compressed afterwards. Popular examples include knowledge distillation Hinton et al. (2015), post-training quantization Jacob et al. (2018), low-rank factorization Hu et al. (2022), and structured pruning Li et al. (2016). All of these methods typically require the costly upfront training of a large network.
In this article, we address the issue of costly pre-training by introducing CompreSSM, a principled in-training compression technique that effectively reduces the dimension of the SSM while largely preserving the expressive power of uncompressed models. We motivate our approach by the control-theoretic origins of SSMs Kalman (1960). In particular, we draw on balanced truncation Antoulas (2005), a classical Model Order Reduction (MOR) technique that approximates a high-dimensional state space system with a low-dimensional one while retaining its essential input–output behavior. Observing that the dominant Hankel Singular Values (HSVs) of an SSM are rank-preserving during training, we propose truncating dimensions associated with small Hankel singular values once they fall below a predefined relative threshold.
@@ 2.1 Discrete Linear Time Invariant Systems
A more general class of systems are Linear Time-Varying (LTV), where the matrices ${\bm{A}},{\bm{B}},{\bm{C}},{\bm{D}}$ are functions of time. Such systems become relevant in the context of selective SSMs, where the system matrices depend on the input. For now, we restrict to the LTI case as the base framework. The LTV case is discussed in Appendix A.2.
@@ 2.2 Spectral stability of Hermitian matrices
Let $\bm{W}$} and $\bm{W}$$^{\prime}$ be Hermitian matrices of size $n$ (i.e. symmetric for real-valued matrices), and let $\delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}$. Also, $\forall i\in[1,\cdots,n]$, let $\lambda_{i}({\bm{W}})$ represent the $i$-th largest eigenvalue of $\bm{W}$}. The ordering $\lambda_{1}({\bm{W}})\geq\cdots\geq\lambda_{n}({\bm{W}})$ can always be established as the eigenvalues of Hermitian matrices are guaranteed to be real.
@@
$\forall i\in[1,\cdots,n]$, $\lambda_{i}({\bm{W}})$ is Lipschitz-continuous on the space of Hermitian matrices with operator norm:
$|\lambda_{i}({\bm{W}}^{\prime})-\lambda_{i}({\bm{W}})|\leq\max_{i=1,\cdots,n}(|\lambda_{i}(\delta{\bm{W}})|)=\max(|\lambda_{1}(\delta{\bm{W}})|,|\lambda_{n}(\delta{\bm{W}})|).$ (7)
@@ 3.1 CompreSSM: the algorithm
Note that the above algorithm can not deliver improved small models if there is no clear correlation between state dimension and model performance. Furthermore, there should be sufficient training in between successive reduction steps for the model to recover from pruning. For more details, we refer the reader to Section 4.2.
The fact that CompreSSM applies an in-training compression scheme enables a significant increase in performance per training time. Indeed, we validate the speedup experimentally in Section 4.2.
@@ 3.2 In-Training Reduction
Figure 2: In-training per-step analysis of Hankel singular value dynamics for a single LRU block with state dimension of 8 on the MNIST dataset for the first 25k steps. The leftmost plot shows the raw HSVs (as a set). The middle-left plot depicts the maximum absolute eigenvalue of $\delta{\bm{H}}$ as described in Section 3.2. The middle-right plot overlays the maximum variation bound as an error margin around each HSV, with each shade now representing a highly probable path for a specific state dimension obtained by step by step linear sum assignment solving. The rightmost plot shows the relative contribution of the bottom $r$ HSVs to the total energy.
@@
We examine the case of a single LRU block trained on the MNIST dataset with state dimension of 8 for visual clarity in Figure 2 (and provide plots for larger models and datasets in Appendix C.2). To keep track of the state dimension to HSV value correspondence, we overlay the maximum absolute eigenvalue of $\delta{\bm{H}}$, on top of each HSV. Empirically, this allows us to robustly identify probable HSV trajectories as the continuity bound is consistently small enough to ensure each eigenvalue is clearly isolated, with minimal bound overlaps and rare gradual HSV relative order crossings occurring (prediction of evolution established via linear sum assignment solution in such cases). Furthermore, the cumulative contribution of the bottom-$r$ HSVs stabilizes rapidly. Indeed, after a small initial number of steps, we observe that both the ordering of singular values remains constant, and dimensions of low importance seldom gain substantial relative energy during training.
@@ 4.1 Experimental Setup
In order to empirically validate in-training balanced truncation, we train a linear recurrent unit (LRU) Orvieto et al. (2023) on datasets of different complexity, ranging from MNIST to tasks from the long range arena (LRA) Tay et al. (2020).
We use the same training pipeline as Rusch & Rus (2025); Walker et al. (2025), accounting for additional quirks of LRU training like a learning rate factor for the sequence mixing layer. The LRA dataloaders are borrowed from Smith et al. (2022), the hyperparameters largely taken from (Orvieto et al., 2023, Table 10). We summarize them in Table 2.
Reduction starts from the full model order reported in column 4 of Table 2. For all datasets but IMDB and MNIST, we attempt four equidistant reduction steps during the warm-up period of the learning rate, which equals $10\%$ of the total steps. Doing so ensures maximal speed up potential for the subsequent $90\%$ of training, while also staying robust to large early training .
As MNIST is trained without learning rate decay, we attempt truncation during all of training. For IMDB, we include a waiting stage and attempt reduction inside a smaller time window. The details can be found in Appendix B.1. Generally, reductions are only executed if the reduced dimension is less than $95\%$ of the current state dimension.
We train non-reduced models with each block initialized at the average final order of the compressed ones to allow for a fair comparison. To increase our baseline statistics and further establish correlations between state dimension and model performance, we train further models with different state dimensions.
@@ 4.2 Empirical Results
However, on datasets where state dimension does clearly correlate with model performance, CompreSSM improves small model performance. On CIFAR10, for example, compressed model performance almost stays constant as a function of the reduction tolerance (and thus for different state dimensions), while the non-compressed counterparts exhibit a approximately $10\%$ performance drop (see entry 1 in Table 1 or Figure 3(a)). The MNIST results paint a similar picture (entry 5 in Table 1 or Figure 4(a)).
@@
The IMDB results reveal the importance of the prerequisites mentioned in Section 3.1 (see Figure 4(d)). Even after significantly reducing the parameters and increasing the droprate (see Appendix B), the unreduced only learn for roughly $8k$ steps before overfitting. However, in order to do in-training balanced truncation, the actual training phase needs to be long enough to allow for a couple of reduction steps with a large enough spacing so that the model can follow the training dynamics for a while without being pruned. Indeed, for non-aggressive pruning (that is, pruning with small tolerance $\tau$), which requires less recovery time, the top reduced models often outperform the baseline.
@@
As mentioned previously, an advantage of CompreSSM is that it comes with a training speedup; as the state dimensions get pruned, training speeds up. Indeed, Figure 3(b) demonstrates this effect on CIFAR10. In the case of state dimension $57$, CompreSSM provides a $7\%$ increase in test accuracy for just $4\%$ more training time.
@@
(b) Test accuracy vs. Training time
Figure 3: Subfigure (3(a)) shows the performance of different models trained on CIFAR10 as a function of the state dimension. Grey data indicates non-reduced models, and the shades of orange correspond to reduced models, with tolerance decreasing with redness. The circles represents the top-3 mean, while the star corresponds to the top-1 model. Subfigure (3(b)) shows top-3 performance versus the normalized average training time. Marker diameter is proportional to the final model order (also annotated) and in-between models are omitted for visual decluttering.
@@ 5 Related Work
Model compression techniques in machine learning. The question of model compression has received considerable attention in the literature (Deng et al., 2020; Zhu et al., 2024), leveraging various techniques (and mixtures thereof) such as pruning (Han et al., 2015; Frantar & Alistarh, 2023), which removes redundant parameters to reduce network size; quantization (Xiao et al., 2022; Gholami et al., 2021), which compresses models by lowering numerical precision; low-rank factorization (Lin et al., 2024), which exploits structure in weight matrices to reduce dimensionality; and knowledge distillation (Hinton et al., 2015; Gou et al., 2020), where a smaller model learns to mimic a larger one.
@@
By contrast, control-theoretic MOR approaches have been applied to diagonal S4 layers, but only as a post-hoc step to initialize retraining (Ezoe & Sato, 2024). Elsewhere, regularizers based on the Hankel nuclear norm or modal $\ell_{1}$ have been shown to encourage parsimonious state representations during training, though without explicit truncation events, and with a price to pay in terms of optimal performance (Forgione et al., 2024). Finally, $\mathcal{H}_{2}$-optimal reductions have been proposed as a competitor to balanced truncation in offline SSM MOR settings (Sakamoto & Sato, 2025).
@@ 6 Conclusion
Finally, we provide an outlook on how to do in-training balanced truncation for linear time-varying systems. In this selective case, we propose averaging the dynamics over the input space and applying the reduction scheme to this time-independent system. For future work, we would like to extend CompreSSM to linear self-attention models, e.g., Gated Linear Attention Yang et al. (2023), Mamba2 Dao & Gu (2024), and Gated DeltaNet Yang et al. (2024), which is another class of models based on linear time-varying systems.
Acknowledgments
This work was supported in part by the Hector Foundation and by the Boeing Company.
Appendix A Appendix

### v3

@@ 1 Introduction
State Space Models (SSMs) (Gu et al., 2021; Hasani et al., 2022; Smith et al., 2022; Orvieto et al., 2023; Rusch and Rus, 2025) have recently emerged as a powerful alternative to established sequence models such as Recurrent Neural Networks (RNNs) and Transformers. They combine the parallelizable training efficiency of scaled dot-product attention (Vaswani, 2017) with the computational and memory advantages of RNNs, enabling strong performance across large-scale language, vision, and audio modeling tasks (Gu and Dao, 2024; Goel et al., 2022; Nguyen et al., 2022).
Despite their efficient structure and recent progress in hardware-aware implementations, current SSMs remain computationally intensive. While both memory and runtime scale with sequence length, the size of the SSM state further amplifies these costs. Reducing the state dimension therefore provides an effective strategy to simultaneously reduce memory usage and runtime. This can be achieved by leveraging techniques from structured compression. However, most existing approaches are commonly applied post-training: a large model is trained to completion and only compressed afterwards. Popular examples include knowledge distillation (Hinton et al., 2015), post-training quantization (Jacob et al., 2018), low-rank factorization (Hu et al., 2022), and structured pruning (Li et al., 2016). All of these methods typically require the costly upfront training of a large network.
In this article, we address the issue of costly pre-training by introducing CompreSSM, a principled in-training compression technique that effectively reduces the dimension of the SSM while largely preserving the expressive power of uncompressed models. We motivate our approach by the control-theoretic origins of SSMs (Kalman, 1960). In particular, we draw on balanced truncation (Antoulas, 2005), a classical Model Order Reduction (MOR) technique that approximates a high-dimensional state space system with a low-dimensional one while retaining its essential input–output behavior. Observing that the dominant Hankel Singular Values (HSVs) of an SSM are rank-preserving during training, we propose truncating dimensions associated with small Hankel singular values once they fall below a predefined relative threshold.
@@ 2.1 Discrete Linear Time Invariant Systems
A more general class of systems are Linear Time-Varying (LTV), where the matrices ${\bm{A}},{\bm{B}},{\bm{C}},{\bm{D}}$ are functions of time. Such systems become relevant in the context of selective SSMs, where the system matrices depend on the input. For now, we restrict to the LTI case as the base framework. The LTV case is discussed in Appendix E.1.
@@ 2.2 Spectral stability of Hermitian matrices
Let $\bm{W}$} and $\bm{W}$$^{\prime}$ be Hermitian matrices of size $n$ (i.e. symmetric for real-valued matrices), and let $\delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}$. Also, $\forall i\in[1,\cdots,n]$, let $\sigma_{i}({\bm{W}})$ represent the $i$-th largest eigenvalue of $\bm{W}$}. The ordering $\sigma_{1}({\bm{W}})\geq\cdots\geq\sigma_{n}({\bm{W}})$ can always be established as the eigenvalues of Hermitian matrices are guaranteed to be real.
@@
$\forall i\in[1,\cdots,n]$, $\sigma_{i}({\bm{W}})$ is Lipschitz-continuous on the space of Hermitian matrices with operator norm:
$|\sigma_{i}({\bm{W}}^{\prime})-\sigma_{i}({\bm{W}})|\leq\max_{i=1,\cdots,n}(|\sigma_{i}(\delta{\bm{W}})|)=\max(|\sigma_{1}(\delta{\bm{W}})|,|\sigma_{n}(\delta{\bm{W}})|).$ (7)
@@ 3.1 CompreSSM: the algorithm
Note that the above algorithm can not deliver improved small models if there is no clear correlation between state dimension and model performance. The fact that CompreSSM applies an in-training compression scheme enables a significant increase in performance per training time. Indeed, we validate the speedup experimentally in Section 4.2, and provide an in-depth complexity and time reduction analysis in Section D of the Appendix.
3.2 Pragmatic variant
In ablations presented in Appendix C, we observe that training can sustain successive moderate balanced truncations without divergence from the large-model upper-bound performance, until impactful HSVs are removed. This observation motivates a slight tweak to the CompreSSM algorithm. Although one does not have access to the upper bound trained at the large dimension throughout, we can track the validation metric during training and assume that, as long as no drop in performance occurs, the reduction does not cause any noticeable global performance drop.
In practice, the procedure is implemented with a simple safeguard. At each fixed-fraction reduction step (10% per reduction is reasonable), we first save the current (pre-reduction) checkpoint. We then apply truncation, train the model for a very small number of steps, and evaluate it on the validation set. If validation performance continues to improve, we proceed to the next reduction. If performance degrades, we discard the reduced version and revert to the previously saved checkpoint, after which no further reductions are applied. Results with this approach are provided in Section B.2.
This protocol ensures that model quality always remains close to that of the unreduced baseline, without the need to explicitly tune the number of reductions or a fixed tolerance level.
@@ 3.3 In-Training Reduction
Figure 2: In-training per-step analysis of Hankel singular value dynamics for a single LRU block with state dimension of 8 on the sMNIST dataset for the first 25k steps. The leftmost plot shows the raw HSVs (as a set). The middle-left plot depicts the maximum absolute eigenvalue of $\delta{\bm{H}}$ as described in Section $3.$ The middle-right plot overlays the maximum variation bound as an error margin around each HSV, with each shade now representing a highly probable path for a specific state dimension obtained by step by step linear sum assignment solving. The rightmost plot shows the relative contribution of the bottom $r$ HSVs to the total energy.
@@
We examine the case of a single LRU block trained on the sMNIST dataset with state dimension of 8 for visual clarity in Figure 2 (and provide plots for larger models and datasets in Appendix B.3). To keep track of the state dimension to HSV value correspondence, we overlay the maximum absolute eigenvalue of $\delta{\bm{H}}$, on top of each HSV. Empirically, this allows us to robustly identify probable HSV trajectories as the continuity bound is consistently small enough to ensure each eigenvalue is clearly isolated, with minimal bound overlaps and rare gradual HSV relative order crossings occurring (prediction of evolution established via linear sum assignment solution in such cases). Furthermore, the cumulative contribution of the bottom-$r$ HSVs stabilizes rapidly. Indeed, after a small initial number of steps, we observe that both the ordering of singular values remains constant, and dimensions of low importance seldom gain substantial relative energy during training.
@@ 4.1 Experimental Setup
In order to empirically validate in-training balanced truncation, we train a linear recurrent unit (LRU) (Orvieto et al., 2023) on datasets of different complexity, ranging from sMNIST to tasks from the long range arena (LRA) (Tay et al., 2020).
We use the same training pipeline as Rusch and Rus (2025); Walker et al. (2025), accounting for additional quirks of LRU training like a learning rate factor for the sequence mixing layer. The LRA dataloaders are borrowed from Smith et al. (2022), the hyperparameters largely taken from (Orvieto et al., 2023, Table 10). We summarize them in Table 3.
Reduction starts from the full model order reported in column 4 of Table 3. For all datasets but IMDB and sMNIST, we attempt four equidistant reduction steps during the warm-up period of the learning rate, which equals $10\%$ of the total steps. Doing so ensures maximal speed up potential for the subsequent $90\%$ of training, while also staying robust to large early training .
As sMNIST is trained without learning rate decay, we attempt truncation during all of training. For IMDB, we include a waiting stage and attempt reduction inside a smaller time window. The details can be found in Appendix A.3. Generally, reductions are only executed if the reduced dimension is less than $95\%$ of the current state dimension.
The baselines we compare against are non-reduced models trained all along with each block initialized at the average final order of the compressed ones to allow for a fair comparison. To increase our baseline statistics and further establish correlations between state dimension and model performance, we train further models with different state dimensions.
@@ 4.2 Empirical Results
However, on datasets where state dimension does correlate with model performance, CompreSSM improves small model performance. On CIFAR10, for example, model performance almost stays constant as a function of reduction tolerance (and thus for different state dimensions), while the non-compressed counterparts exhibit a approximately $10\%$ performance drop (see entry 1 in Table 1 or Figure 3(a)). The sMNIST results paint a similar picture (entry 5 in Table 1 or Figure 4(a)).
@@
The IMDB results reveal the importance of the prerequisites mentioned in Section 3.1 (see Figure 4(d)). Even after significantly reducing the parameters and increasing the droprate (see Appendix A), the unreduced only learn for roughly $8k$ steps before overfitting. However, in order to do in-training balanced truncation, the actual training phase needs to be long enough to allow for a couple of reduction steps with a large enough spacing so that the model can follow the training dynamics for a while without being pruned. Indeed, for non-aggressive pruning (that is, pruning with small tolerance $\tau$), the top reduced models often outperform the baseline.
@@
As mentioned previously, an advantage of CompreSSM is that it comes with a training speedup; as the state dimensions get pruned, training speeds up. Indeed, Figure 3(b) demonstrates this effect on CIFAR10. CompreSSM at dimension $92$ preserves nearly full baseline accuracy ($85.7\%$ vs. $86.5\%$) and achieves a $1.5\times$ speedup. By comparison, training directly at dimension $92$ is only marginally faster ($1.6\times$ speedup) yet markedly worse, reaching just $81.8\%$ accuracy. A detailed breakdown of in-training speedup gains is provided in Appendix D.
@@
(b) Test accuracy vs. Training speedup factor
Figure 3: Subfigure (3(a)) shows the performance of different models trained on CIFAR10 as a function of the state dimension. Grey data indicates non-reduced models, and the shades of orange correspond to reduced models, with tolerance decreasing with redness. The circles represents the top-3 mean, while the star corresponds to the top-1 model. Subfigure (3(b)) shows top-3 performance as a function of training speedup with respect to the original large model training without reductions ($n=384$ baseline). Marker diameter is proportional to the final model order (also annotated) and in-between models are omitted for visual decluttering.
4.3 Comparison to Other Model Order Reduction Techniques
We compare CompreSSM to other model order reduction approaches, namely reduction with Hankel Nuclear Norm (HNN) regularization (Forgione et al., 2024; Schwerdtner et al., 2025), as well as Knowledge Distillation (KD) (Hinton et al., 2015). The approaches are described in Appendix F, with implementation details provided as well. Results of the experiments are summarized in Table 2.
The former experiment reveals three clear findings: (i) HNN regularization is computationally expensive, slowing training by an order of magnitude at least due to repeated Gramian-eigenvalue evaluations (at every single gradient step); (ii) HNN-constrained models underperform unconstrained baselines even before reduction; and (iii) when targeting small final dimensions, CompreSSM achieves higher accuracy while being over dramatically faster.
On the other hand, knowledge distillation performs on par with CompreSSM as long as the teacher has a similar state dimension as the student. However, as the model is further compressed, KD’s performance drops significantly, while CompreSSM is able to maintain performance. KD furthermore requires training the large teacher model to completion before training a smaller student model, as well as doing a forward pass through the teacher even when training the student. This can be substantially slower than CompreSSM.
Table 2: MOR techniques on sMNIST and CIFAR10. Top-3 mean test accuracies (%) and training speedup factors are reported. HNN Regularization is evaluated on sMNIST, and Knowledge Distillation is evaluated on CIFAR10. $n$ represents models’ final order (rounded).
Method Metric $\mathbf{n=13}$ $\mathbf{n=28}$ $\mathbf{n=47}$ $\mathbf{n=76}$ $\mathbf{n=148}$ $\mathbf{n=191}$ $\mathbf{n=256}$
sMNIST Baseline Acc. (%) $92.6$ $96.0$ $95.9$ $96.4$ $\mathbf{97.3}$ $\mathbf{97.3}$ $\mathbf{97.3}$
Speed $\times$ $3.1$ $2.8$ $2.7$ $2.4$ $2.0$ $1.7$ $1.0$
CompreSSM Acc. (%) $\mathbf{95.9}$ $\mathbf{96.9}$ $\mathbf{96.9}$ $\mathbf{96.9}$ $97.0$ $97.2$
Speed $\times$ $2.8$ $2.6$ $2.5$ $2.3$ $1.9$ $1.6$
HNN Reg. Acc. (%) $91.7$ $95.8$ $95.8$ $95.8$ $95.8$ $95.8$ $95.9$
Speed $\times$ $0.06$ $0.06$ $0.06$ $0.06$ $0.06$ $0.06$ $0.06$
Method Metric $\mathbf{n=57}$ $\mathbf{n=93}$ $\mathbf{n=126}$ $\mathbf{n=161}$ $\mathbf{n=214}$ $\mathbf{n=327}$ $\mathbf{n=384}$
CIFAR10 Baseline Acc. (%) $78.2$ $81.8$ $83.7$ $84.2$ $84.9$ $86.0$ $86.5$
Speed $\times$ $1.69$ $1.62$ $1.53$ $1.43$ $1.22$ $1.03$ $1.00$
CompreSSM Acc. (%) $\mathbf{84.4}$ $\mathbf{85.7}$ $\mathbf{86.0}$ $\mathbf{85.8}$ $\mathbf{86.0}$ $86.1$
Speed $\times$ $1.58$ $1.52$ $1.41$ $1.33$ $1.17$ $1.03$
KD Acc. (%) $79.4$ $83.5$ $84.4$ $85.3$ $\mathbf{86.0}$ $\mathbf{87.0}$
Speed $\times$ $0.55$ $0.52$ $0.61$ $0.51$ $0.49$ $0.45$
@@ 5 Related Work
Model compression techniques in machine learning. The question of model compression has received considerable attention in the literature (Deng et al., 2020; Zhu et al., 2024), leveraging various techniques (and mixtures thereof) such as pruning (Han et al., 2015; Frantar and Alistarh, 2023), which removes redundant parameters to reduce network size; quantization (Xiao et al., 2022; Gholami et al., 2021), which compresses models by lowering numerical precision; low-rank factorization (Lin et al., 2024), which exploits structure in weight matrices to reduce dimensionality; and knowledge distillation (Hinton et al., 2015; Gou et al., 2020), where a smaller model learns to mimic a larger one.
@@
By contrast, control-theoretic MOR approaches have been applied to diagonal S4 layers, but only as a post-hoc step to initialize retraining (Ezoe and Sato, 2024). Elsewhere, regularizers based on the Hankel nuclear norm or modal $\ell_{1}$ have been shown to encourage parsimonious state representations during training, though without explicit intermediate truncation events, and with a price to pay in terms of both per step compute time as well as optimal performance (Forgione et al., 2024; Schwerdtner et al., 2025). Finally, $\mathcal{H}_{2}$-optimal reductions have been proposed as a competitor to balanced truncation in offline SSM MOR settings (Sakamoto and Sato, 2025). This approach is computationally expensive, requiring gradient on a complex optimization problem, without the solution guaranteeing stability of the reduced order system.
Although not a compression approach, the authors in Yu et al. (2024) approach the problem of Hankel singular value efficiency from the perspective of SSM initialization, providing an approach to avoid fast decaying singular values.
@@ 6 Conclusion
Finally, we provide an outlook on how to do in-training balanced truncation for linear time-varying systems. In this selective case, we propose averaging the dynamics over the input space and applying the reduction scheme to this time-independent system. For future work, we would like to extend CompreSSM to linear self-attention models, e.g., Gated Linear Attention (Yang et al., 2023), Mamba2 (Dao and Gu, 2024), and Gated DeltaNet (Yang et al., 2024), which is another class of models based on linear time-varying systems.
Appendix A Implementation Details